OpenStudy (prominer10000):

PLEASE HELP! WILL MEDAL & FAN! Which is an equation in point-slope form for the given point and slope? Point: (8, -4) ; Slope: -5/6 A. y+4=-5/6(x+8) B. y-4=-5/6(x+8) C. y+4=-5/6(x-8) D. y-4=5/6(x-8)

OpenStudy (will.h):

we will have to use the following form y - y1 = m(x - x1) m is the slope and we have the point (8,-4) where y1 is -4 and x1 is 8 So substitute in the original form y - (-4) = -5/6 (x - 8) simplify the left side y + 4 = -5/6(x - 8)

OpenStudy (prominer10000):

Oh ok, so it's C?

OpenStudy (will.h):

indeed

OpenStudy (prominer10000):

Ok thanks! :)